# Low-Latency ML Inference by Grouping Correlated Data Objects and Computation

### v1

@@ 3.3. Potential Benefits
A platform-level runtime engine can take advantage of correlation keys in many different ways.
@@
Proactive collocation: It is possible to proactively collocate correlated data and computation by factoring correlation keys into the caching infrastructure, as well as into placement decisions of both computations on servers and objects within a scalable storage infrastructure. For example, in the RCP application, all positions from the same agent (output from the MOT step) can be stored in the same physical node where the PRED computation for that agent will take place when/if it arrives. In this example, the correlation key for all data requests and for the computation request should be the same (the identifier of the agent, for example). Collocation avoids the extra overhead of fetching input data from remote processes. Section 4 shows evidence of this benefit: as seen in Figure 5, end-to-end latency is significantly lower and more consistent when employing correlation grouping.
Prefetching: The knowledge encoded into correlation keys can also enable proactive decisions in anticipation of a future need. For example, in a multistage job a scheduler launching the first stage of a computation might also dispatch a notification that some collection of objects will be required in the second or subsequent stages. This is why our correlation key mechanism permits one object to be tagged with multiple keys: if we consider an ML computation such as RCP with a graphical structure, the output of one stage might be an input to multiple downstream stages, each with its own group of data dependencies. Using correlation keys, this would enable the platform to prefetch objects that will be needed but are not currently in cache, to transfer objects from originating node to a consuming node in anticipation of the future need, etc. Notice the similarity here to a publish-subscribe model, in which publishers might not know who the subscribers will be, and subscribers might not know their publishers. In our work, the correlation key is playing an analogous role.
@@
Consistency: Many edge ML applications are sensitive to data sequencing. Similarly to partition keys in Azure EH, correlation keys can be used to guarantee the consistency of requests. Requests with the same correlation key may have to be handled sequentially and in order. In platforms where objects sharing the same correlation key are to be stored on the same “home nodes”, one update can be used to update the entire set as a single atomic action.
@@
Parallelism: Requests with different correlation keys have no dependency with each other and may thus be handled in parallel. Furthermore, the correlation grouping mechanism enables a finer grain of parallelism control compared, for example, with a mechanism such as Azure EHs. In EHs, requests with different partition keys may be placed in a same partition, causing the requests to be handled sequentially, although they could be potentially have been parallelized.
4. Scaling with Correlation Grouping
@@
In this section, we describe an implementation of the correlation grouping mechanism in Cascade (Song et al., 2022; Derecho Project, 2023): a state-of-the-art stream processing platform targeted at latency-sensitive ML workflows. Cascade was selected because it offered the highest baseline performance among open source platforms amenable to our methodology, enabling us to ask whether even lower latency and higher throughput might be feasible using an correlation grouping methodology.
@@
The goal of this section is to show the benefit of correlation grouping on the end-to-end (E2E) latency of the RCP application. We evaluate the E2E latency of the application with different workloads, while also scaling out the Cascade deployment. Section 5 provides a deeper analysis on the overheads in each step of the pipeline.
@@ 4.1. RCP Application Implementation
Our experiment used the Stanford Drone Dataset (SDD) (Robicquet et al., 2016), which contains 60 aerial videos showing campus intersections. The videos include several types of agents (pedestrians, cars, cyclists, skateboarders, buses, and golf carts). Data source “clients” stream these videos at 2.5 FPS. All three steps are implemented in Python using the PyTorch library.
To implement the MOT step, a multi-object tracker (Broström, 2022) that employs YOLO5 for agent detection was used. StrongSORT (Du et al., 2022) and OSNet (Zhou et al., 2022b) are used for re-identification and trajectory tracking. We selected three videos from the SDD for training (little3, hyang5, and gates3). Data is uncompressed, hence each frame is approximately 8MB in size. State data containing positions and features of agents detected in a frame (required to re-identify agents in the following frame) ranges from a few KBs to around 10MB depending on the number of agents in the frame. Frames from the same video must be processed sequentially, but frames from different videos can be processed in parallel by multiple instances of the MOT step. Appendix A includes plots showing the number of agents detected in each frame of each video, as well as the size of the MOT state.
The PRED step employs a trajectory prediction model, YNet (Mangalam et al., 2021) pretrained on the SDD data by the authors (Girase, 2022). PRED requires eight consecutive positions for each agent and makes no prediction if fewer than eight are available. Position objects are small: 10s of bytes. Instances of PRED for the same agent must be processed sequentially. Instances triggered for different agents may be executed in parallel. Notice that the PRED workload depends not only on how many clients are streaming frames, but also on how many agents are detected in each frame.
CD consists of a simple algorithm that performs a linear interpolation on the predicted trajectories of identified agents, outputting a warning if any pair crosses. Instances of this step corresponding to the same frame and client must be processed sequentially to ensure that all pairs of trajectories will be matched with each other. Different frames (even if from the same client) may be processed in parallel.
@@ 4.2. Cascade: System Architecture
Cascade is a full-stack platform for high-speed stream processing that prioritizes low latency by hosting data and compute, avoiding copying or locking on critical data paths and leveraging acceleration technologies such as RDMA or DPDK. Backend components are implemented in C++, while client APIs are available in multiple languages (e.g. Java and Python). Cascade consists of a set of nodes interconnected in a complete network. Nodes can be clients or servers. Client nodes may issue requests to any server node. Server nodes implement two subsystems: storage and computation. Each subsystem is further described next. Figure 2 shows an overview of the system architecture.
Figure 2. Cascade architecture.
@@
Cascade’s storage subsystem implements a sharded Key/Value (K/V) object store. Server nodes are logically grouped into disjoint shards with a few nodes each. Prior to our work, each object key was hashed to determine the home shard for the K/V pair. Cascade supports multiples levels of persistence, but our work considered only the two in-memory mode: trigger requests (which cause a task to run but leave no data behind) volatile storage requests (which replicate an object that will then be retained in memory by all members of the home shard). Volatile storage is limited to objects that can fit within memory. In the RCP scenario, for example, object sizes ranged from a few hundred bytes to a few MBs.
Cascade’s User-Defined Logic (UDL) framework is responsible for executing computational tasks. The code for each task is supplied by the developer as a container or in a DLL. In either case, a self-initialization method is expected to register one or more upcallable functions, each associated with a key prefix. On each Cascade server, an upcall will occur if that server receives a K/V pair with a matching key. For example, if a task is registered using the prefix “/RCP/taxis”, then an update to a K/V pair with key “/RCP/taxis/1234” would trigger that task. Cascade’s trigger operation will cause a single upcall to occur, on whichever node the put was sent to. Cascade’s volatile put has two possible behaviors, under developer control: it can upcall on every node in the shard that the key mapped to, or it can do the upcall on one randomly selected node within the shard.
Although a single K/V trigger or volatile put will trigger a task, once the computation is running that task can access additional objects using Cascade’s K/V get API. Moreover, any task can issue a trigger put or a volatile put, enabling pipelined tasks in which each stage initiates the next stage.
The potential overheads in this model are thus: (1) a P2P send operation is used to send a trigger put; (2) an atomic multicast is used to send a volatile put; (3) if a task calls get to access an object that is not homed on the node where it is executing, it must either be fetched from cache or over the network. Even with high-speed networking primitives supported by RDMA or DPDK, network transfers of large objects are costly, so Cascade must minimize the frequency of such events. Nodes have limited capacity, hence a further issue arises if the execution framework needs to enlarge the pool of compute resources. Lacking a warm-start mechanism, Cascade will incur the cost of fetching all needed objects.
@@
Cascade supports a form of resource partitioning called an object pool. The basic idea is as follows: rather than treating all Cascade nodes as a single sharded service, the nodes are instead organized into groups. A single server node can belong to multiple pools. Cascade allows the developer to configure each pool with its own properties, such as the shard size to use, and the degree of data persistence the pool will offer. A persistent pool can be pinned to specific storage hardware (SSD, rotating disk, Optane, etc). A volatile pool can be pinned to host memory, GPU memory, etc. Pools are identified by pathname prefixes: if /x/y is the prefix of a volatile pool holding GPU memory objects, /x/y/z could name a tensor residing within that GPU memory.
4.3. Correlation Grouping on Cascade
@@
Most aspects of the Cascade architecture can easily accommodate the proposed correlation grouping mechanism. As noted earlier, data and computation requests in Cascade consist of trigger, put or get operations. The first two are parameterized by the key and value (uninterpreted byte vector) of the object being transmitted/saved, whereas get takes a key and returns the corresponding object. We implemented function $f$ as a regular expression, which is matched against request keys, but with one simplification: our initial implementation limits each request to a single correlation key  [^1] In future work we will support mappings to multiple correlation keys. A single key suffices for the RCP application, but multiple keys are likely to be important in more complex applications where the output of one compute stage is consumed by multiple follow-on stages, each with its own dependency set.. The correlation key is a substring of the request key, composed of the characters that matched the regular expression. We employed the Hyperscan library (Wang et al., 2019) for matching regular expressions, which introduced a negligible overhead to Cascade’s critical path: according to microbenchmarks, matching the regular expressions employed in the RCP application was under 300 microseconds on average.
@@
To implement the application-level API of the correlation grouping mechanism, we extended the Cascade client API. The extension allows developers to register regular expressions with a specified object pool. Figure 3 shows a code excerpt (in Python) with an example of how object pools are created, with and without employing correlation grouping. The only modification made to the Cascade API was the addition of the optional argument correlation_set_regex to the method that creates an object pool (line 12). Any put or get operation assigns an correlation key to the corresponding object by matching the object key against the regular expression registered in the corresponding object pool (line 18).
@@
5# creating object pool without correlation grouping
@@
9# creating object pool with correlation grouping
@@
14# putting object in the object pool that is not grouped
@@
17# putting object in the object pool that is grouped
18capi.put("/grouping/example_1",None) # correlation key is ’_1’
Figure 3. Code excerpt showing how object pools with/without grouping are created and how objects are put/get.
With respect to the platform-level runtime engine, we made two modifications to Cascade: (i) the mapping from key to shard within an object pool is based on the correlation key instead of the request key; and (ii) the correlation key functions and their matching expressions are registered in all nodes. As described above, Cascade previously selected the shard that will handle a request by hashing the object key supplied with the object describing the request. Our modified policy selects the shard by hashing the correlation key. Jointly, these changes ensure that all objects with the same correlation key will be stored (and replicated) in the same shard, and route requests using that same correlation key to this shard.
An open question concerns whether (and how) schedulers should be extended to factor in correlation grouping considerations. Cascade is already designed to send computations to the home nodes of needed objects, or to nodes where those objects are cached: an ideal policy from our perspective.
@@ 4.4. Computing Environment
We now describe a series of experiments conducted on our local cluster. Servers in the cluster are equipped with Mellanox ConnectX-4 VPI NIC cards connected to a Mellanox SB7700 InfiniBand switch. This configuration results in a RDMA-capable 100Gbps network backbone. All servers have their clocks synchronized using PTP (Eidson et al., 2002), making timestamps from different servers comparable with sub-millisecond precision. Servers have two configurations, denoted A and B in this work. Configuration A consists of two Intel Xeon Gold 6242 CPUs, 192 GB of memory, and an NVIDIA Tesla T4 GPU. This configuration is commonplace in Edge Cloud environments, where applications such as RCP are often deployed. Configuration B consists of two Intel Xeon E2690 v0 CPUs and 96 GB of memory, but no GPU. All servers run Linux kernel 5.15.0. Cascade nodes were deployed on Docker (version 24.0.6) containers hosted on the cluster servers. Containers were based on Ubuntu 22.04, with CUDA version 11.7.
@@
Cascade was deployed with three object pools, one per step. Cascade nodes in pools responsible for MOT and PRED were deployed on containers hosted on servers with configuration A, since these steps require GPUs for ML inference. Clients and other Cascade nodes were deployed on containers hosted on servers with configuration B. In this section, we report results with a varying number of shards in each pool, as well as a varying number of nodes in each shard. Each configuration of shards/nodes is called a layout.
@@
Although Cascade includes a scheduler, we configured the system to place objects and trigger computations purely by correlation key hashing. This enables us to focus on the degree to which a purely correlation grouping placement of data and computation can improve latency.
4.5. RCP Application Deployment on Cascade
@@
Each step of the RCP application was deployed as a DLL on Cascade’s UDL framework. UDLs are loaded by each Cascade node on startup. ML models are also loaded on startup by the UDLs – all nodes have local access to models and hyperparameters. All required objects (MOT state data, agent positions and trajectory predictions) are stored/retrieved using Cascade K/V store.
@@
Each client is responsible for a different video, and sends complete video frames at a rate of 2.5 FPS. To send a frame, a client puts an object in the K/V store with a key that will trigger a MOT task. The MOT task retrieves the state data from the previous frame, and puts a separate object for each agent position found in the frame. Each agent position put by the MOT task triggers a PRED computation. Each PRED task first retrieves the past positions of the corresponding agent from the K/V store, and puts an object with the trajectory prediction. Each trajectory prediction triggers a CD computation, which first retrieves all trajectory predictions for the same frame available in the K/V store, and then performs the computation. The result of the collision detection is put in the K/V store. The E2E latency is the time elapsed between the client sending the frame, until the result of the last collision detection for that frame is put in the K/V store. All tasks cache in memory all objects they retrieve or create.
@@
We employed two different workloads: single clients (little3, hyang5, and gates3), and three simultaneous clients (little3 + hyang5 + gates3). Each client send 700 frames, and we discard measurements from the first 100 frames. Each experiment is repeated three times. Cascade was completely cleared of objects and object pools before each run of each experiment, and a setup script was executed to create the necessary object pools.
@@
We grouped requests from each step using the correlation grouping mechanism in a different way. For MOT, grouping was based on the identifier of the client, so that all frames from a same client always went to the same shard. For PRED, grouping was based on agent identifiers, and for CD it was based on both the client identifier and the frame number. Table 1 shows all the employed object pools, examples of object keys, which computations are triggered when objects are put in the corresponding object pool (if any), the regular expressions used to group requests (if any), and the resulting correlation keys from matching the regular expressions against the example keys.
@@
Table 1. Object pools, example keys, triggered steps, regular expressions, and correlation keys.
Object Pool Example Key Step Regex Correlation Key
/collision/tracking/cameras /collision/tracking/cameras/little3_42 MOT /[a-zA-Z0-9]+_ /little3_
/collision/tracking/states /collision/tracking/states/little3_42 /[a-zA-Z0-9]+_ /little3_
/collision/tracking/agent_position /collision/tracking/agent_position/little3_7_42 PRED /[a-zA-Z0-9]+_[0-9]+_ /little3_7_
/collision/prediction/agent_prediction /collision/prediction/agent_prediction/little3_42_7 CD /[a-zA-Z0-9]+_[0-9]+_ /little3_42_
/collision/detection /collision/detection/little3_42_7_5
@@
Results reported in this section compare two different placement strategies, random placement and correlation grouping. In the random placement strategy, the RCP application is executed without using correlation grouping, i.e. objects and computations are placed randomly by hashing requests keys – the standard Cascade behavior. In the correlation grouping strategy, we group requests as described above.

### v2

@@ 3.4. Potential Benefits
A platform-level runtime engine can take advantage of affinity keys in many different ways.
@@
Proactive collocation: It is possible to proactively collocate correlated data and computation by factoring affinity keys into the caching infrastructure, as well as into placement decisions of both computations on servers and objects within a scalable storage infrastructure. For example, in the RCP application, all positions from the same actor (output from the MOT step) can be stored in the same physical node where the PRED computation for that actor will take place when/if it arrives. In this example, the affinity key for all objects and tasks should be the same (the identifier of the actor, for example). Collocation avoids the extra overhead of fetching input data from remote processes.
Prefetching: The knowledge encoded into affinity keys can also enable proactive decisions in anticipation of a future need. For example, in a multistage job a scheduler launching the first stage may decide to prefetch objects that will be needed by a downstream stage, based on the affinity key of the objects and the corresponding stage, in anticipation of the computational task reaching the corresponding nodes.
@@
Consistency: Many edge AI applications are sensitive to data sequencing. Similarly to partition keys in Azure EH, affinity keys can be used to guarantee consistency. Objects and tasks with the same affinity key may have to be handled sequentially and in order. Additionally, a platform may allow objects sharing the same affinity key to be updated atomically, in a single action: this was easily achieved in our prototype, since objects with the same affinity key are stored in a same shard, as described later in Section 4.
@@
Parallelism: Objects and tasks with different affinity keys have no mutual dependency and may thus be handled in parallel. Furthermore, the affinity grouping mechanism enables a finer grain of parallelism control compared, for example, with a mechanism such as Azure EHs. In EHs, requests with different partition keys may be placed in the same partition, causing the requests to be handled sequentially.
4. Evaluation on Local Cluster
@@
In this section, we describe an implementation of the affinity grouping mechanism in Cascade (Song et al., 2023): a state-of-the-art stream processing platform targeted at latency-sensitive AI workflows. Cascade was selected because it offered the highest baseline performance among open source platforms amenable to our methodology, enabling us to ask whether even lower latency and higher throughput might be feasible using an affinity grouping methodology.
@@
The goal of this section is to show the benefit of affinity grouping on the end-to-end (E2E) latency of the RCP application. We evaluate the E2E latency of the application with different workloads, while also scaling out the Cascade deployment. Section 5 provides a deeper analysis on the overheads in each step of the pipeline.
@@ 4.1. RCP Application Implementation
Our experiment used the Stanford Drone Dataset (SDD) (Robicquet et al., 2016), which contains aerial videos showing campus intersections. The videos include several types of actors (e.g. pedestrians, cars, cyclists). Data source “clients” simulate cameras by streaming these videos to the pipeline. The MOT and PRED steps are based on off-the-shelf AI models, and we used the source code and pre-trained models provided by the authors as much as possible: most of our modifications focused on data ingress/egress. All three steps are implemented in Python using the PyTorch library.
To implement the MOT step, a multi-object tracker (Broström, 2022) that employs YOLO5 for actor detection was used. StrongSORT (Du et al., 2022) and OSNet (Zhou et al., 2022b) are used for re-identification and trajectory tracking. We selected three videos from the SDD for training (little3, hyang5, and gates3). Data is uncompressed, hence each frame is approximately 8MB in size (compressing and then decompressing frames is slower than just transferring them uncompressed). State data containing positions and features of actors detected in a frame (required to re-identify actors in the following frame) ranges from a few KBs to around 10MB depending on the number of actors in the frame. Frames from the same video must be processed sequentially, but frames from different videos can be processed in parallel by multiple MOT instances.
The PRED step employs a trajectory prediction model, YNet (Mangalam et al., 2021) pretrained on the SDD data by the authors (Girase, 2022). PRED requires eight consecutive positions for each actor and makes no prediction if fewer than eight are available. Position objects are small: 10s of bytes. Instances of PRED for the same actor must be processed sequentially. Instances triggered for different actors may be executed in parallel. Notice that the PRED workload depends not only on how many clients are streaming frames, but also on how many actors are detected in each frame by MOT.
CD consists of a simple algorithm that performs a linear interpolation on the predicted trajectories of identified actors, outputting a warning if any pair crosses. Instances of this step corresponding to the same frame and client must be processed sequentially to ensure that all pairs of trajectories will be matched with each other. Different frames (even if from the same client) may be processed in parallel.
@@ 4.2. Cascade: System Architecture
Cascade (Song et al., 2023) is a full-stack platform for high-speed stream processing that prioritizes low latency by hosting data and compute, avoiding copying and locking on critical data paths and leveraging acceleration technologies such as RDMA and DPDK. It consists of a set of nodes (clients and servers) interconnected in a complete network, typically in an edge/cloud datacenter. Client nodes may issue requests to any server node. Server nodes implement two subsystems: storage and computation. Each subsystem is further described next. Figure 2 shows an overview of the system architecture.
@@
The storage subsystem implements a sharded Key/Value (K/V) object store. Server nodes are logically grouped into disjoint shards. Prior to our work, each object key was hashed to determine the home shard for the K/V pair. Cascade supports multiples levels of persistence, but our work considered only the two in-memory mode: trigger requests, which cause a task to run but leave no data behind; and volatile storage requests, which replicate an object that will then be retained in memory by all members of the home shard.
Cascade’s User-Defined Logic (UDL) framework is responsible for executing computational tasks. The code for each task is supplied by the developer as a container or in a Dynamic-Link Library (DLL). In either case, a task is associated with a key prefix. On each Cascade server, an upcall will occur if that server receives a K/V pair with a matching key. For example, if a task is registered using the prefix “/RCP/taxis”, then an update to a K/V pair with key “/RCP/taxis/1234” would trigger that task at whichever node the put was sent to. Each task can also issue another put, enabling pipelined tasks in which each stage initiates the next stage. In the context of this work, a potential overhead that we want to minimize occurs when a task reads an object that is not homed on the node where it is executing. The object then must either be fetched from cache or over the network. Even with high-speed networking primitives supported by RDMA or DPDK, network transfers of large objects are costly.
@@
Cascade supports a form of resource partitioning called an object pool. Rather than treating all Cascade nodes as a single sharded service, the nodes are instead organized into groups. A single server node can belong to multiple pools. Cascade allows the developer to configure each pool with its own properties, such as the shard size to use, and the degree of data persistence the pool will offer. For example, a volatile pool can be pinned to host memory, GPU memory, etc. Pools are identified by pathname prefixes: if /x/y is the prefix of a volatile pool holding GPU memory objects, /x/y/z could name a tensor residing within that GPU memory.
4.3. Affinity Grouping on Cascade
@@
As noted earlier, requests in Cascade consist of trigger, put or get operations. The first two are parameterized by the key and value (uninterpreted byte vector) of the object being transmitted/saved, whereas get takes a key and returns the corresponding object. We implemented function $f$ as a regular expression, which is matched against request keys (the descriptor). The affinity key is a substring of the request key, composed of the characters that matched the regular expression. We employed the Hyperscan library (Wang et al., 2019) for matching regular expressions, which introduced a negligible overhead to Cascade’s critical path: according to microbenchmarks, matching the regular expressions employed in the RCP application was under 300 microseconds on average.
@@
To implement the application-level API of the affinity grouping mechanism, we extended the Cascade client API. The extension allows developers to register regular expressions with a specified object pool. Listing 1 shows a code excerpt (in Python) with an example of how object pools are created, with and without employing affinity grouping. The only modification made to the Cascade API was the addition of the optional argument affinity_set_regex to the method that creates an object pool (line 12). Any put or get operation assigns an affinity key to the corresponding object by matching the object key against the regular expression registered in the corresponding object pool (line 18).
Listing 1: Creating pools with/without affinity grouping.
@@
5# creating object pool without affinity grouping
@@
9# creating object pool with affinity grouping
@@
14# putting in the object pool that is not grouped
@@
17# putting in the object pool that is grouped
18capi.put("/grouping/example_1",None) # affinity key is ’_1’
With respect to the platform-level runtime engine, we made two modifications to Cascade: (i) the mapping from key to shard within an object pool is based on the affinity key instead of the request key; and (ii) the affinity key functions and their matching expressions are registered in all nodes. As described above, Cascade previously selected the shard that will handle a request by hashing the object key supplied with the object describing the request. Our modified policy selects the shard by hashing the affinity key. Jointly, these changes ensure that all objects with the same affinity key will be stored (and replicated) in the same shard, and route requests using that same affinity key to this shard.
@@ 4.4. Computing Environment
Servers in our local cluster are equipped with Mellanox ConnectX-4 VPI NIC cards connected to a Mellanox SB7700 InfiniBand switch, resulting in a RDMA-capable 100Gbps network backbone. All servers have their clocks synchronized using PTP (Eidson et al., 2002), making timestamps from different servers comparable with sub-millisecond precision. Servers have two configurations, denoted A and B in this work. Configuration A consists of two Intel Xeon Gold 6242 CPUs, 192 GB of memory, and an NVIDIA Tesla T4 GPU. Configuration B consists of two Intel Xeon E2690 v0 CPUs and 96 GB of memory, but no GPU. Our experiments employ up to 8 servers with configuration A, and up to 9 with B.
@@
Cascade was deployed with three object pools, one per step. Cascade nodes in pools responsible for MOT and PRED were deployed on servers with configuration A, since these steps require GPUs for AI inference. Clients and other Cascade nodes were deployed on servers B. In this section, we report results with a varying number of shards in each pool. Each shard has always only one node (i.e. a physical server): increasing the size of each shard would increase replication, which is not relevant for evaluating our proposals. Each configuration of shards is called a layout, and we denote a layout as $x/y/z$, where values $x$, $y$, and $z$ indicate the number of shards for steps MOT, PRED, and CD, respectively.
@@
Although Cascade includes a scheduler, we configured the system to place objects and trigger computations purely by affinity key hashing. This enables us to focus on the degree to which a purely affinity grouping placement of data and computation can improve latency.
4.5. RCP Deployment on Cascade
@@
Each step of the RCP application was deployed as a DLL on Cascade’s UDL framework. UDLs are loaded by each Cascade node on startup, along with models, weights, and hyperparameters. All required objects (MOT state data, actor positions and trajectory predictions) are stored/retrieved using Cascade K/V store.
@@
Each client is responsible for a different video, and sends video frames at a rate of 2.5 FPS (the rate the AI model employed in PRED was trained at). To send a frame, a client puts an object in the K/V store with a key that will trigger a MOT task. The MOT task retrieves the state data from the previous frame, and puts a separate object for each actor position found in the frame. Each actor position put by the MOT task triggers a separate PRED computation. Each PRED task first retrieves the past positions of the corresponding actor from the K/V store, and puts an object with the trajectory prediction. Each trajectory prediction triggers a separate CD computation, which first retrieves all trajectory predictions for the same frame available in the K/V store, and then performs the computation. The result of the collision detection is put in the K/V store. The E2E latency is the time elapsed between the client sending the frame, until the result of the last collision detection for that frame is put in the K/V store. All tasks cache in memory all objects they retrieve or create.
@@
We employed two workloads: single clients (little3, hyang5, and gates3), and three simultaneous clients (little3 + hyang5 + gates3). Each client sends 700 frames, and we discard measurements from the first 100 frames. Each experiment is repeated three times. Cascade was completely cleared of objects and object pools before each run of each experiment.
@@
We grouped requests from each step using the affinity grouping mechanism in a different way. For MOT, grouping was based on the identifier of the client, so that all frames from the same client always went to the same shard. For PRED, grouping was based on actor identifiers, and for CD it was based on both the client identifier and the frame number. Table 1 shows all the employed object pools, key examples, triggered tasks (if any), the regular expressions used to group requests (if any), and the resulting affinity keys.
@@
Table 1. Object pools, example keys, triggered steps, regular expressions, and affinity keys.
Object Pool Example Key Step Regex Affinity Key
/frames /frames/little3_42 MOT /[a-zA-Z0-9]+_ /little3_
/states /states/little3_42 /[a-zA-Z0-9]+_ /little3_
/positions /positions/little3_7_42 PRED /[a-zA-Z0-9]+_[0-9]+_ /little3_7_
/predictions /predictions/little3_42_7 CD /[a-zA-Z0-9]+_[0-9]+_ /little3_42_
/cd /cd/little3_42_7_5
@@
Results reported in this section compare two different placement strategies, random placement and affinity grouping. In the random placement strategy, the RCP application is executed without using affinity grouping, i.e. objects and computations are placed randomly by hashing requests keys – the standard Cascade behavior. In affinity grouping, we group requests as described above. Hash-based pseudo-random mapping from key to shard is standard in key-value stores. Our discovery is that by combining affinity grouping with randomization on the affinity key, we get the best of both worlds: excellent load-balancing and scaling with sharply higher cache hit rates, hence less data movement.